OpenStudy (anonymous):

How to find x in cos2x = 3/4? Please explain in detail.

OpenStudy (anonymous):

cos2x=2cos^2 x-1=3/4 2cos^2 x=3/4+1=7/4 cos^2 x=7/8 taking square root cosx=0.93541 x=20.705

OpenStudy (anonymous):

Yes, that's the value I get too. But, the question gives the interval 0 <= x <= pi. How does this work?

OpenStudy (swissgirl):

Is it \( cos^2(x) \) or \(Cos(2x)\) ?

OpenStudy (anonymous):

It is cos(2x).

OpenStudy (anonymous):

\[2\cos^2(x)-1=\frac{3}{4}\] may help

OpenStudy (anonymous):

well not really i think you can say \(2x=\cos^{-1}(\frac{3}{4})\) whatever that is, then divide by 2

OpenStudy (anonymous):

Sorry - I should have included. In the memo, the answers are given: "x = 0.36 and x = 2.78."

OpenStudy (anonymous):

???

OpenStudy (swissgirl):

Not sure how that can be the answer

OpenStudy (swissgirl):

I got x=20.70481105

OpenStudy (anonymous):

I'll give the full question, but I'm not sure how it'll help.

OpenStudy (swissgirl):

This is a diff question

OpenStudy (swissgirl):

a=3 and b=-4 So the equation would be f(x)=3-4Cos(2x)

OpenStudy (anonymous):

No, go through the steps. I found that a = 3 and b = -4 at (i). Then, plugging those in into f(x), simplifies to cos2x = 3/4. This is where I'm stuck.

OpenStudy (swissgirl):

ohhhhhhhh i see

OpenStudy (anonymous):

Do you have any idea where to go from here? Thanks for helping so far, by the way!

OpenStudy (swissgirl):

Okkk I got both answers. Umm What calculator are you using?

OpenStudy (anonymous):

I haven't got it with me here, but it's a fully adequate scientific calculator. Why? :-)

OpenStudy (swissgirl):

ok Well I was getting the wrong answer since my calculator was on degrees setting when it should basically when you solve this on your calc make sure its on radians

OpenStudy (anonymous):

Really? So how did you get those answers?

OpenStudy (swissgirl):

\(2\cos^2(x)-1=\large\frac{3}{4}\) \(2\cos^2(x)=1.75\) \(\cos^2(x)=.875\) \(\cos(x)=\sqrt{.875} \) \(\cos(x)= \pm .9354143467\) \(x=\cos^{-1}( .9354143467) \text{ and } x=\cos^{-1}( .9354143467)\) \(x=.3613671239, and, x=2.78022553\)

OpenStudy (anonymous):

Alright! I understand everything from the first line downwards, but you have me confused right at the top. How does \[\cos2x = \frac{ 3 }{ 4 }\] become \[2\cos^2(x) - 1 = \frac{ 3 }{ 4 }?\] I know the rule \[\cos^2x + \sin^2x = 1\] but I'm not sure how it applies here. You're a gem for helping me, thank you so much!

OpenStudy (swissgirl):

Its a trig identity \(\cos(2x) = \cos^2(x) – \sin^2(x) = 1 – 2\sin^2(x) = 2\cos^2(x) – 1\)

OpenStudy (anonymous):

I didn't know that one! I'll remember it now. Thank you SO much for your help. It feels silly just giving you a 'best response' but it's all I can do. I hope you don't mind if I ask though, but why do you do this? It's so kind and selfless, it's really great of you. Thanks again!

OpenStudy (swissgirl):

lol I used to ask loads of questions on here too so I know what you mean. I like answering questions just to refresh my memory and its kinda fun answering questions that you know the answer to.

OpenStudy (anonymous):

Well, you've saved my life.

OpenStudy (anonymous):

By the way, if I close this question, will I be able to come back to it?

OpenStudy (swissgirl):

Yes on your top right hand side you will see your username in blue Click on that and a new page will open Then on your left hand side you will see Questions Asked in blue. Click on that and you will be able to see all the questions you have asked

OpenStudy (anonymous):

Thanks! Another thing, sorry to bother, but that trig identity, does it apply to sin and tan as well?

OpenStudy (swissgirl):

sin(2x) = 2sin(x)cos(x) \(tan(2x) = \large{2 tan(x) \over 1 - tan^2(x)}\)

OpenStudy (swissgirl):

http://www.purplemath.com/modules/idents.htm I am getting all my info from here. I dont know all the trig identities off by heart.

OpenStudy (anonymous):

I think that's all I have to ask for now. Goodnight :-)

OpenStudy (swissgirl):

Gnite :)